# thanks: Corresponding author

### v1

@@ Appendix B Constructive counterexample in Sch-AdS demonstrating that $r_{\rm ps}(r_{h})$ can decrease
where $M$ is the ADM mass and $P=-\Lambda/(8\pi)$ denotes the thermodynamic pressure. The photon-sphere condition $rf^{\prime}(r)-2f(r)=0$ gives
@@
Figure 8: Plots of $r_{\rm ps}(r_{h})$ for the Sch-AdS black hole. Colored solid curves denote nine different pressures $P$, and the black dashed curve shows the path-constrained case with $k<0$ and $b>0$.
@@
Appendix C Local Existence of $M(r_{h})$ and $P(r_{h})$
@@
Appendix D Proof of the Positivity of $\partial_{P}\Phi_{1}$
@@
$\mathrm{sgn}\,\partial_{P}\Phi_{1}=\begin{cases}+,&\Delta_{x}<0\ \text{or}\ x\in(0,x_{-})\cup(x_{+},\infty),\\[2.0pt] 0,&x=x_{\pm},\\[2.0pt] -,&x\in(x_{-},x_{+}).\end{cases}$ (105)
@@
Fixing $M$, $q$, and $d$ at finite values and taking the limit $P\to\infty$, the condition $\Phi_{1}(r,M,q,P,d)=0$ can be satisfied only if
@@
This shows that the quantities $x_{\pm}^{1/(d-3)}$ correspond exactly to the limiting values of $r_{\rm isco}$ as $P\to\infty$. Next, we examine the limit $P\to 0$. For $d=4$, one obtains
@@
  • $q\neq 0$: For the RN black hole, it is well known that $r_{\mathrm{isco}}$ attains its minimum at $M=q$, where $r_{\mathrm{isco,min}}=4M$. From Eq. (104), $x_{+}$ decreases monotonically with $q$, reaching its maximum $x_{+,{\rm max}}={15M}/{4}$ at $q=0$, so that $x_{min}\!\mid_{P=0}=4M>x_{+,{\rm max}}={15M}/{4}$.
@@
Next, consider increasing $P$ from zero in Eq. (112). Initially, one has $\Phi_{1}=\widetilde{\Phi}_{1}(r_{\mathrm{isco}}\!\mid_{P=0},M,q,d)=0$ with $\Gamma_{1}=\partial_{r}\widetilde{\Phi}_{1}>0$. Once $P$ increases, the term $\partial_{P}\Phi_{1}>0$ appears, so $\widetilde{\Phi}_{1}$ must decrease to maintain $\Phi_{1}=0$. Since $\Gamma_{1}>0$, this requires $r_{\mathrm{isco}}$ to decrease, meaning that $x$ moves toward $x_{+}$. As $P$ continues to increase, we assume continuous dependence of $r_{\mathrm{isco}}$ and $\partial_{P}\Phi_{1}$ on $P$, so that $x$ must vary continuously from $x\!\mid_{P=0}$ to $x_{+}\!\mid_{P\to\infty}$.
@@
For higher dimensions ($d>4$), the argument becomes even simpler. At $P=0$, Schwarzschild and RN black holes possess no ISCO ($r_{\mathrm{isco}}\to\infty$). As $P$ increases, $r_{\mathrm{isco}}$ moves inward from infinity toward $x_{+}$, again ensuring $\partial_{P}\Phi_{1}>0$. Hence, in all cases we conclude that
@@
Appendix E Asymptotic Relations in the Small-Horizon Limit
@@
For $\partial_{P}\Phi_{1}$, since $r_{h}\to 0^{+}$ implies $P\to\infty$, one has $\partial_{P}\Phi_{1}\to 0$ (see Appendix D for details). The thermodynamic volume takes the simple form
@@
For the ISCO, the divergence $P\!\to\!\infty$ implies (see Appendix D)
@@
Appendix F Monotonicity of $r_{\rm c}(r_{h})$ in the Phase-Transition Region
F.1 Photon sphere ($\epsilon=0$)
Our proof consists of two steps. First, we demonstrate that $r_{h0}$ lies to the left of $r_{h,\min}$. Second, we show that $P(r_{h0})>P(r_{h,\max})$, where $r_{h,\min}$ and $r_{h,\max}$ denote the locations of the minimum and maximum of $P(r_{h})$, respectively. This procedure circumvents the need to determine the small-black-hole radius $r_{h,s}$ from Maxwell’s equal-area law while still guaranteeing that $r_{h0}$ lies to the left of $r_{h,s}$.
Step 1
From Eq. (41) it follows that ${\mathrm{d}P}/{\mathrm{d}r_{h}}=0$ implies ${\mathrm{d}M}/{\mathrm{d}r_{h}}>0$. Since $M(r_{h})$ first decreases and then increases, $r_{h0}$ must therefore lie to the left of both extrema, namely on the small-black-hole branch.
Step 2
To establish the result for arbitrary $q$, we eliminate $q$ by expressing $P(r_{h})$ and ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ in reduced variables:
$P(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{1-6\tilde{r}_{h}^{2}+8\tilde{r}_{h}^{3}\tilde{T}_{0}}{3\tilde{r}_{h}^{4}},$ (125)
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{-1+3\tilde{r}_{h}^{2}+4\tilde{r}_{h}^{3}\tilde{T}_{0}}{6\tilde{r}_{h}^{2}},$ (126)
where $\tilde{r}_{h}=r_{h}/r_{h,c}$ and $\tilde{T}_{0}=T_{0}/T_{c}$, with $r_{h,c}=\sqrt{6}\,q$ and $T_{c}=1/(3\sqrt{6}\,\pi q)$ denoting the critical values. Using these relations, we numerically determine $r_{h0}$ and $r_{h,\max}$, and plot $P(r_{h0})$ (red curve) and $P(r_{h,\max})$ (blue curve) as functions of $\tilde{T}$ in Fig. 9. The black dashed vertical line at $\tilde{T}_{0}=\sqrt{2}/2$ marks the critical value above which $P>0$; the region $\tilde{T}_{0}<\sqrt{2}/2$ is therefore unphysical and excluded from consideration. From the remaining portion of the plot it is evident that $P(r_{h0})>P(r_{h,\max})$. Combined with the result of Step 1, this demonstrates that $r_{h0}$ lies to the left of $r_{h,s}$.
To visualize this behavior, Fig. 10 shows $P(r_{h})$ (red) and ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ (blue) for several $\tilde{T}_{0}$. At $\tilde{T}_{0}=0.4$, $P(r_{h0})<P(r_{h,\max})$, but $P(r_{h})<0$ and is thus unphysical. When $\tilde{T}_{0}\geq\sqrt{2}/2$, $P(r_{h})\geq 0$ and $P(r_{h0})>P(r_{h,\max})$, with $r_{h,s}$ and $r_{h,l}$ clearly located to the right of $r_{h0}$.
F.2 ISCO ($\epsilon=1$)
With the photon-sphere case established, the ISCO analysis becomes straightforward. From the criterion

### v2

@@ Appendix B Constructive counterexample in Sch–AdS demonstrating that $r_{\rm ps}(r_{h})$ can decrease
where $M$ is the ADM mass and $P=-\Lambda/(8\pi)$ denotes the thermodynamic pressure. The photon sphere condition $rf^{\prime}(r)-2f(r)=0$ gives
@@
Figure 8: Plots of $r_{\rm ps}(r_{h})$ for the Sch–AdS black hole. Colored solid curves denote nine different pressures $P$, and the black dashed curve shows the path-constrained case with $k<0$ and $b>0$.
@@
Appendix C Conditions of the photon sphere and ISCO
For photons ($\epsilon=0$) the radial function (5) reduces to
$\mathcal{R}(r)=E^{2}-\frac{f(r)\,L^{2}}{r^{2}}.$ (88)
Imposing the circular-orbit condition, $\mathcal{R}^{\prime}(r_{c})=0$, gives
$\mathcal{R}^{\prime}(r_{\rm ps})=-\frac{f^{\prime}(r_{\rm ps})\,L^{2}}{r_{\rm ps}^{2}}+\frac{2f(r_{\rm ps})\,L^{2}}{r_{\rm ps}^{3}}=0.$ (89)
Eliminating $L^{2}$ yields
$\Phi_{0}(r_{\rm ps}):=r_{\rm ps}\,f^{\prime}(r_{\rm ps})-2f(r_{\rm ps})=0,$ (90)
which corresponds to the photon-sphere equation (7). Imposing the instability of the photon circular orbit, $\mathcal{R}^{\prime\prime}(r_{\rm ps})>0$, gives
$\mathcal{R}^{\prime\prime}(r_{\rm ps})=-\frac{L^{2}}{r_{\rm ps}^{3}}\,\underbrace{\big[r_{\rm ps}f^{\prime\prime}(r_{\rm ps})-f^{\prime}(r_{\rm ps})\big]}_{\;\equiv\ \partial_{r}\Phi_{0}(r_{\rm ps})}>0.$ (91)
Since $L^{2}/r_{\rm ps}^{3}>0$, it follows that
$\Gamma_{0}:=\partial_{r}\Phi_{0}(r_{\rm ps})=r_{\rm ps}f^{\prime\prime}(r_{\rm ps})-f^{\prime}(r_{\rm ps})<0,$ (92)
which corresponds to the stability criterion (9).
For massive particles ($\epsilon=1$) the radial function (5) reduces to
$\mathcal{R}(r)=E^{2}-f(r)\!\left(1+\frac{L^{2}}{r^{2}}\right).$ (93)
The conditions
$\mathcal{R}^{\prime}(r_{\rm isco})=\mathcal{R}^{\prime\prime}(r_{\rm isco})=0,$ (94)
give two algebraic equations for $L$ and $r_{\rm isco}$:
$\mathcal{R}^{\prime}(r_{\rm isco})=\frac{2f(r_{\rm isco})\,L^{2}}{r_{\rm isco}^{3}}-f^{\prime}(r_{\rm isco})\!\left(1+\frac{L^{2}}{r_{\rm isco}^{2}}\right)=0,$ (95)
$\mathcal{R}^{\prime\prime}(r_{\rm isco})=-\frac{6f(r_{\rm isco})L^{2}}{r_{\rm isco}^{4}}$ $+\frac{4f^{\prime}(r_{\rm isco})L^{2}}{r_{\rm isco}^{3}}$ (96)
$-f^{\prime\prime}(r_{\rm isco})\!\left(1+\frac{L^{2}}{r_{\rm isco}^{2}}\right)=0.$
Eliminating $L^{2}$ yields
$\mathcal{R}^{\prime\prime}(r_{\rm isco})=\frac{2\,\Phi_{1}(r_{\rm isco})}{r_{\rm isco}^{2}\,f^{\prime}(r_{\rm isco})-2\,r_{\rm isco}f(r_{\rm isco})}=0,$ (97)
where we define
$\Phi_{1}(r_{\rm isco})=r_{\rm isco}\,f(r_{\rm isco})\,f^{\prime\prime}(r_{\rm isco})$ $-2\,r_{\rm isco}\,[f^{\prime}(r_{\rm isco})]^{2}$ (98)
$+3\,f(r_{\rm isco})\,f^{\prime}(r_{\rm isco})=0$
which corresponds to the ISCO equation (11). To determine the sign of
$\Gamma_{1}:=\partial_{r}\Phi_{1}(r_{\rm isco}),$ (99)
we rewrite Eq. (97) as
$\mathcal{R}^{\prime\prime}(r_{\rm c})=\frac{2\,\Phi_{1}(r_{\rm c})}{\,r_{\rm c}^{2}f^{\prime}(r_{\rm c})-2r_{\rm c}f(r_{\rm c})\,}$, (100)
where $r_{\rm c}$ denotes the radius of a timelike circular orbit in the neighborhood of the ISCO. By definition of the ISCO, orbits with $r_{\rm c}>r_{\rm isco}$ are stable $\bigl(\mathcal{R}^{\prime\prime}(r_{\rm c})<0\bigr)$, whereas those with $r_{\rm c}<r_{\rm isco}$ are unstable $\bigl(\mathcal{R}^{\prime\prime}(r_{\rm c})>0\bigr)$. Next, We analyze the sign of $\mathcal{R}^{\prime\prime}(r_{\rm c})$, which is fixed by the numerator and denominator of the ratio above. We abbreviate the denominator as
$D(r_{\rm c}):=r_{\rm c}^{2}f^{\prime}(r_{\rm c})-2r_{\rm c}f(r_{\rm c})=r_{\rm c}^{4}\,\partial_{r}\!\left(\frac{f}{r^{2}}\right)\Big|_{r_{\rm c}},$ (101)
where $V(r)\propto f(r)/r^{2}$ is the photon effective potential. For $r>r_{\rm ps}$ one has $\partial_{r}V<0$, and circular orbits near the ISCO satisfy $r_{\rm c}>r_{\rm ps}$. Therefore $D(r_{\rm c})<0$. We now examine the numerator. Expanding $\Phi_{1}$ about the ISCO,
$\Phi_{1}(r_{\rm c})=\Phi_{1}(r_{\rm isco})+\partial_{r}\Phi_{1}(r_{\rm isco})\,(r_{\rm c}-r_{\rm isco})+O\bigl[(r_{\rm c}-r_{\rm isco})^{2}\bigr],$ (102)
and using $\Phi_{1}(r_{\rm isco})=0$, we obtain near $r_{\rm isco}$:
$\mathrm{sgn}\left[\mathcal{R}^{\prime\prime}(r_{\rm c})\right]=\mathrm{sgn}\!\left[\frac{\partial_{r}\Phi_{1}(r_{\rm isco})\,(r_{\rm c}-r_{\rm isco})}{D(r_{\rm isco})}\right].$ (103)
Consequently, to ensure $\mathcal{R}^{\prime\prime}(r_{\rm c})>0$ for $r_{\rm c}<r_{\rm isco}$ and $\mathcal{R}^{\prime\prime}(r_{\rm c})<0$ for $r_{\rm c}>r_{\rm isco}$, the slope of $\Phi_{1}$ at the ISCO must be positive, $\Gamma_{1}:=\partial_{r}\Phi_{1}(r_{\rm isco})>0$, which is precisely the stability criterion (12).
Appendix D Local Existence of $M(r_{h})$ and $P(r_{h})$
@@
Appendix E Proof of the Positivity of $\partial_{P}\Phi_{1}$
@@
$\mathrm{sgn}\left(\partial_{P}\Phi_{1}\right)=\begin{cases}+,&\Delta_{x}<0\ \text{or}\ x\in(0,x_{-})\cup(x_{+},+\infty),\\[2.0pt] 0,&x=x_{\pm},\\[2.0pt] -,&x\in(x_{-},x_{+}).\end{cases}$ (114)
@@
Fixing $M$, $q$, and $d$ at finite values and taking the limit $P\to+\infty$, the condition $\Phi_{1}(r,M,q,P,d)=0$ can be satisfied only if
@@
This shows that the quantities $x_{\pm}^{1/(d-3)}$ correspond exactly to the limiting values of $r_{\rm isco}$ as $P\to+\infty$. Next, we examine the limit $P\to 0$. For $d=4$, one obtains
@@
  • $q\neq 0$: For the RN black hole, it is well known that $r_{\mathrm{isco}}$ attains its minimum at $M=q$, where $r_{\mathrm{isco,min}}=4M$. From Eq. (113), $x_{+}$ decreases monotonically with $q$, reaching its maximum $x_{+,{\rm max}}={15M}/{4}$ at $q=0$, so that $x_{\rm min}\!\mid_{P=0}=4M>x_{+,{\rm max}}={15M}/{4}$.
@@
Next, consider increasing $P$ from zero in Eq. (121). Initially, one has $\Phi_{1}=\widetilde{\Phi}_{1}(r_{\mathrm{isco}}\!\mid_{P=0},M,q,d)=0$ with $\Gamma_{1}=\partial_{r}\widetilde{\Phi}_{1}>0$. Once $P$ increases, the term $\partial_{P}\Phi_{1}>0$ appears, so $\widetilde{\Phi}_{1}$ must decrease to maintain $\Phi_{1}=0$. Since $\Gamma_{1}>0$, this requires $r_{\mathrm{isco}}$ to decrease, meaning that $x$ moves toward $x_{+}$. As $P$ continues to increase, we assume continuous dependence of $r_{\mathrm{isco}}$ and $\partial_{P}\Phi_{1}$ on $P$, so that $x$ must vary continuously from $x\!\mid_{P=0}$ to $x_{+}\!\mid_{P\to+\infty}$.
@@
For higher dimensions ($d>4$), the argument becomes even simpler. At $P=0$, Schwarzschild and RN black holes possess no ISCO ($r_{\mathrm{isco}}\to+\infty$). As $P$ increases, $r_{\mathrm{isco}}$ moves inward from infinity toward $x_{+}$, again ensuring $\partial_{P}\Phi_{1}>0$. Hence, in all cases we conclude that
@@
The limit used above,
$\lim_{P\to 0}r_{\rm isco}=+\infty,$ (125)
can be argued as follows. For the higher-dimensional Schwarzschild black hole ($q=0$, $d>4$), the ISCO condition takes the form
$\Phi_{1}(r_{\rm isco})$ $=\frac{16\pi P\,r_{\rm isco}}{(d-2)(d-1)}\Bigl[8-c_{d}(d^{2}-1)M\,r_{\rm isco}^{3-d}\Bigr]$ (126)
$\quad-c_{d}(d-3)M\,r_{\rm isco}^{2(1-d)}\Bigl[\,c_{d}(d-1)M\,r_{\rm isco}^{3}$
$\qquad\qquad\qquad\qquad\qquad+(d-5)\,r_{\rm isco}^{d}\Bigr].$
If $r_{\rm isco}$ remains finite as $P\to 0$, the first (pressure-dependent) term vanishes and $\Phi_{1}$ reduces to
$\Phi_{1}(r_{\rm isco})=-c_{d}(d-3)M\,r_{\rm isco}^{2(1-d)}$ $\Bigl[\,c_{d}(d-1)M\,r_{\rm isco}^{3}$ (127)
$+(d-5)\,r_{\rm isco}^{d}\Bigr],$
which admits no positive root for $r_{\rm isco}$ when $d>4$. Thus a finite $r_{\rm isco}$ is incompatible with the ISCO condition in this limit. By contrast, if $r_{\rm isco}\to+\infty$ as $P\to 0$, the second term in $\Phi_{1}$ is suppressed, and one finds
$\Phi_{1}(r_{\rm isco})\;\simeq\;\frac{128\pi P\,r_{\rm isco}}{(d-2)(d-1)},$ (128)
so that the product $P\,r_{\rm isco}$ can tend to zero as $P\to 0$. Therefore the ISCO condition can be satisfied only if
$\lim_{P\to 0}r_{\rm isco}=+\infty.$ (129)
For the higher-dimensional RN black hole ($q\neq 0$, $d>4$), the ISCO condition takes the form
$\Phi_{1}(r_{\rm isco})=$ $-2(d-3)\,q\,r_{\rm isco}^{\,2(1-2d)}\bigl(-q\,r_{\rm isco}^{3}+r_{\rm isco}^{d}\bigr)^{2}$ (130)
$\times\bigl[2(d-2)\,q\,r_{\rm isco}^{3}+(d-5)\,r_{\rm isco}^{d}\bigr]$
$+\frac{32\pi P\,r_{\rm isco}}{(d-2)(d-1)}\Bigl\{4-(d-1)\,q\,r_{\rm isco}^{\,3-2d}$
$\qquad\times\bigl[-2(d-2)\,q\,r_{\rm isco}^{3}+(d+1)\,r_{\rm isco}^{d}\bigr]\Bigr\},$
where we have imposed the extremality condition $M=2q/c_{d}$. Repeating the same analysis as in the Schwarzschild case then yields Eq. (125).
Appendix F Asymptotic Relations in the Small-Horizon Limit
@@
For $\partial_{P}\Phi_{1}$, since $r_{h}\to 0^{+}$ implies $P\to+\infty$, one has $\partial_{P}\Phi_{1}\to 0$ (see Appendix E for details). The thermodynamic volume takes the simple form
@@
For the ISCO, the divergence $P\!\to\!+\infty$ implies (see Appendix E)
@@
Appendix G Monotonicity of $r_{\rm c}(r_{h})$ in the Phase Transition Region
We take the four-dimensional case as an example and show that the phase transition occurs only along the increasing branch of $r_{\rm c}(r_{h})$. The proof strategy is simple: we numerically scan the parameter space, extract the extremal radius $r_{h,e}$ of $r_{\rm ps}(r_{h})$ and the small–black hole radius $r_{h,s}$ at phase coexistence, and then compare their values. In the isothermal ensemble the free parameters are the temperature $T_{0}$ and the charge $q$. Conveniently, the explicit qq–dependence can be removed by introducing reduced (dimensionless) variables
$\tilde{r}_{h}=\frac{r_{h}}{r_{h,c}},\quad\tilde{M}=\frac{M}{M_{c}},\quad\tilde{P}=\frac{P}{P_{c}},\quad\tilde{T}=\frac{T}{T_{c}},$ (140)
where the critical values at the thermodynamic critical point are
$r_{h,c}$ $=\sqrt{6}\,Q,\qquad M_{c}=\frac{2\sqrt{6}\,Q}{3},$ (141)
$P_{c}$ $=\frac{1}{96\pi\,Q^{2}},\qquad T_{c}=\frac{1}{3\sqrt{6}\,\pi\,Q}.$
They are obtained by imposing the criticality conditions
$\frac{\partial P}{\partial r_{h}}=\frac{\partial^{2}P}{\partial r_{h}^{2}}=0,\qquad\frac{\partial T}{\partial r_{h}}=\frac{\partial^{2}T}{\partial r_{h}^{2}}=0.$ (142)
Using these reduced variables, we further obtain
$\tilde{P}(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{1-6\tilde{r}_{h}^{2}+8\tilde{r}_{h}^{3}\tilde{T}_{0}}{3\tilde{r}_{h}^{4}},$ (143)
$\frac{\mathrm{d}\tilde{M}}{\mathrm{d}\tilde{r}_{h}}(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{-1+3\tilde{r}_{h}^{2}+4\tilde{r}_{h}^{3}\tilde{T}_{0}}{6\tilde{r}_{h}^{2}},$ (144)
which no longer contain $q$ explicitly, so the conclusions derived from them apply to arbitrary charge. The only remaining control parameter is $\tilde{T}_{0}$. To ensure $\tilde{P}(\tilde{r}_{h},\tilde{T}_{0})\geq 0$ for all $\tilde{r}_{h}$, one must have $\tilde{T}_{0}\geq\sqrt{2}/2$, while the first–order phase transition occurs only for $T_{0}<T_{c}$, i.e., $\tilde{T}_{0}<1$. Thus the relevant range is $\tilde{T}_{0}\in[\sqrt{2}/2,\,1)$.
Within this interval, we numerically extract $r_{h,s}$ from $\tilde{P}(\tilde{r}_{h},\tilde{T}_{0})$ via the Maxwell equal-area construction, and obtain $r_{h,e}$ from the zero of $\mathrm{d}\tilde{M}/\mathrm{d}\tilde{r}_{h}$. The resulting curves are shown in Fig. 9. From this figure, one clearly sees that $r_{h,e}<r_{h,s}$ for all $\tilde{T}_{0}$ in the allowed range. Equivalently, the zero of $\mathrm{d}r_{\rm ps}/\mathrm{d}r_{h}$ always lies to the left of $r_{h,s}$, demonstrating that the phase transition takes place on the increasing branch of $r_{\rm ps}(r_{h})$.
We now extend this conclusion to the ISCO case. From the criterion (38), i.e.,